The 3s orbital has two radial nodes. Radial nodes occur at Rn,l(r) = 0. Determine the values r for those nodes in terms of a0. (Hint: Focus on the quadratic equation in the parentheses only and solve it for ρ. Thus the roots or the two values are yielded for ρ. Continue on to solve for r twice using ρ = (2Z/na0)r.)
Answer #1

a Radial ware function for as-arbital Rus · Il p2-6P+6) 2-82 913 for R3s P? = 0, 6P + Ő = 0. r Qx2+bx+c 7 Bett L X=-bI56²-4ac